What volume of water should be added to 3.23g of copper (II) chloride (CuCl2) to form a 0.100 mol dm-3 solution?


1
Expert's answer
2021-11-09T10:24:37-0500

Molarity=(g/li)/RFM

g/li=0.1*134.45

g/li=13.445

If 13.445g=1000ml

3.23g=?

(3.23*1000)/13.445

=240.24ml